OpenStudy (anonymous):

If g(x) = x2 + 3, find g(4). 11 19 16 8 I Got B @texaschic101

OpenStudy (anonymous):

x^2 typo

OpenStudy (texaschic101):

g(x) = x^2 + 3 g(4) = 4^2 + 3 g(4) = 16 + 3 g(4) = 19 yep..your right again

OpenStudy (anonymous):

Thanks

OpenStudy (texaschic101):

YW :)